OpenStudy (gracie):

how do u 7c3d2/21c2d3

OpenStudy (anonymous):

=c/3d

OpenStudy (katrinakaif):

Okay lets break this apart: First take the first term which is 7/21 = 1/3 Then next: c^3/c^2 = c3-c2 = c^1 Then Last: d^2/d^3 = d2-d3 = d^-1 which means that we have to put d^-1 in the denominator as it is negative Gving our final answer to be 1c/3d

OpenStudy (anonymous):

7c^3d^2/21c^2d^3 Let's take this term by term. 7 * c^3 * d^2 / 21 * c^2 * d^3 The 7/21 becomes 1/3 The c^3/c^2 is the law of exponents, so you have: c^3-2 = c^1 = c The d^2/d^3 is the law of exponents, so you have: d^2-3 = d^-1 = 1/d So your final answer is: 1/3 * c * 1/d which can be rewritten as: c/3d If this helps, please choose good answer. Visit my website at: http://www.tutorsean.net for free math videos.